Future value of ordinary annuity


Bambi is planning for her retirement. She deposits 300,000 at the end of each year into an account paying 5% interest compounded annually

(a) how much would the account be worth after 10 years?

(b) how much would the account be worth after 20 years?

(c) when Bambi retires in 30 years , what will be the total worth of the account?

(d) If Bambi found a bank that paid 65 interest compounded annually, rather than 5%,how much more would she have in the account after 30 years.


1
Expert's answer
2021-03-25T03:18:10-0400

"FVOA\ =A\ *\ \frac{{(1+r)}^{n-1}}{r}"


"(a) \ \ FVOA\ =300,000\ *\ \frac{{(1+0.05)}^{10}-1}{0.05}\ =\ 3,773,367.76"


"(b)\ \ FVOA\ =300,000\ *\ \frac{{(1+0.05)}^{20}-1}{0.05}\ =\ 9,919,786.23"


"(c)\ \ FVOA\ =300,000\ *\ \frac{{(1+0.05)}^{30}-1}{\ 0.05}\ \ =\ 19,931,654.25"


"(d)\ \ FVOA\ =300,000\ *\ \frac{{(1+0.65)}^{30}-1}{0.65}\ =\ 1,544,280,169,864.45"


"bambi\ would\ earn\ 1,544,260,238,210.20\ more\ after\ 30years"